# Detailed nuclear structure calculations for coherent elastic neutrino-nucleus scattering

### v1

@@ IV.1 Evaluating the nuclear responses
where $J_{i}$ is the nuclear ground state angular momentum, and the nuclear operators $X=M,\,\,\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are given in Appendix C.
The nature of these operators can be understood from their long-wavelength ($q\rightarrow 0$) limit (provided in Table 3). The $M$ operator is spin-independent (SI) and adds coherently from all nucleons, giving rise to the standard coherent enhancement of CEvNS. The $\Phi^{\prime\prime}$ term in the $q\rightarrow 0$ limit is proportional to a spin-orbit operator, and is sensitive to nucleon occupations of the spin-orbit partner levels. This partially coherent operator is highest when one spin-orbit partner is filled and the other is empty. The $\Sigma^{\prime}$ operator is spin-dependent (SD), picking out the transverse projection of the spin current. All nuclei have non-zero responses to the $M$ and $\Phi^{\prime\prime}$ operators, while only odd-nuclei with a $J_{i}\neq 0$ ground state have a $\Sigma^{\prime}$ response.
@@
Using the Wigner-Eckhart theorem the nuclear matrix elements in Eq. (IV.1) can be reduced in angular momentum and isospin:
@@
where $\tilde{a}_{\beta}=(-1)^{j_{\beta}-m_{j_{\beta}}+1/2-m_{t_{\beta}}}\ a_{|\beta|;-m_{j_{\beta}},-m_{t_{\beta}}}$ and $\otimes$ denotes a tensor product. This basis gives response functions of the form $e^{-2y}p(y)$, where $p(y)$ is a polynomial with $y=(qb/2)^{2}$ and $b={1}/{\sqrt{m_{N}\omega}}\approx\sqrt{41.467/(45A^{-1/3}-25A^{-2/3})}$ fm is the harmonic oscillator length parameter ($\omega$ is the oscillator frequency).
IV.2 Form factors from shell model calculations
Large-scale nuclear shell model calculations were performed using NuShellX [37], which is often employed to obtain theoretical energies, wave functions and spectroscopic overlaps for a wide range of nuclei.The program’s inputs include the relevant valence space for the nucleus/isotope of interest - which constitutes a set of single nucleon levels above a completely filled nuclear core - as well as its corresponding list of two-body interactions. Each interaction is fitted to a set of experimental nuclear data for a range of nuclei in the relevant valence space. The contemporary approach is to derive the interactions from the nucleon-nucleon interaction using effective field theory. The interactions are then fine tuned by fitting a limited number of parameters to a body of nuclear data - usually excitation energies in a range of nuclei in the valence space. The aforementioned OBDMEs are extracted for the valence space levels, with the values dependent on the particular shell model interactions used.
In this work we explore a variety of shell model interactions and truncations for our chosen targets: 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs. Using the shell model calculations detailed below, we extract the OBDME for all targets. The OBDMEs are then used to compute the nuclear responses using the Mathematica package described in [38]. Analytic form factor expressions for 40Ar, 70,72,73,74,76Ge and 127I are provided in [39], while those for 133Cs are provided in Appendix D. The form factors for each of the nuclear responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are plotted as functions of $q$ in Appendix D. The form factors are used to derive point-proton, point-neutron and weak radii for all isotopes for comparison with those obtained with our SHF calculations, and are given in Table 7 of Appendix A. Additionally, we make comparisons with experimental energy spectra, moments and transitions in Appendix B.
The shell-model derived Weak form factors are shown in Fig. 2, compared to form factors based on the charge distribution and SHF calculations. The uncertainty band denotes the range of values obtained given the different shell model interactions. The shell model results for each isotope show good agreement with each other (low overall uncertainty) and fall within the large uncertainty band of the form factors derived from the charge distribution. They are also generally in very good agreement with the SHF results within the experimentally accessible regions of interest. The axial form factor is shown in Fig. 3 for each interaction individually.
The sections below outline the valence space, valence space truncation and shell model interactions used in this work for each target.
@@ IV.2 Form factors from shell model calculations
Figure 3: Spin-dependent form factors relevant to CEvNS for the three $J_{i}\neq 0$ targets we consider. The curves denote the different shell model interactions as described in the text.
40Ar
The $sdpf$ valence space was employed for this nucleus, which constitutes single particle levels $1d_{5/2}$, $2s_{1/2}$, $1d_{3/2}$, $1f_{7/2}$, $2p_{3/2}$, $1f_{5/2}$ and $2p_{1/2}$. The valence space truncation employed consisted of an unrestricted $sd$ shell and an empty $pf$ shell for the protons, and a full $sd$ shell with an unrestricted $pf$ shell for the neutrons. Four shell model interactions were used, namely SDPF-NR [40], SDPF-U [41], EPQQM [42] and SDPF-MU [43], all of which consider a 16O core. This isotope was not considered in the work of [35, 38].
The 40Ar theoretical energy levels for the above shell model interactions are plotted against the experimental values in Fig. 5. The first two theoretical energy levels for all interactions are consistent with experiment within $\sim 200$ keV, except for SDPF-MU which is within $\sim 400$ keV from experiment for the $2_{1}^{+}$ state. Overall, the energy spectrum of the EPQQM interaction is least consistent with the other shell model calculations and with experiment, with the spectrum being much more sparse and spread out in the level spacings.
The magnetic dipole and electric quadrupole moments and transitions are presented in Table 8, where the effective $g$ factors and charges employed take on values of $e_{p}=1.5$, $e_{n}=0.5$, $g_{sp}=5.586$, $g_{sn}=-3.826$, $g_{lp}=1$, and $g_{ln}=0$. Information on the ground state is most important for comparison with theoretical values, as it is used in the scattering calculations here. However, as the ground state of 40Ar is $0^{+}$, the magnetic dipole and electric quadrupole moment values are identically zero. Nevertheless, we employ the next non-zero state ($2_{1}^{+}$), which can give an indication of the overall accuracy of the calculation through information on the spectrum beyond the ground state. It is found that the shell model moments are not consistent with experimental values, and most experimental transitions are poorly reproduced as well.
The obtained form factors are compared against those based on the SHF calculations and the charge distribution in Appendix D. The maximum difference between the form factors for the $M$ response at the minimum $q$ value is $\sim 0.3\%$, and $\sim 11\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 24\%$ and $\sim 21\%$, respectively, where the largest differences occur between the SDPF-MU and EPQQM interactions.
70,72,73,74,76Ge
The germanium shell model calculations were performed in the $f_{5}pg_{9}$ model space, which consists of the single-particle levels $2p_{3/2}$, $1f_{5/2}$, $2p_{1/2}$ and $1g_{9/2}$. The work of [35, 38] employed the GCN2850 interaction [45], with a valence space truncation that consisted of limiting the occupation number of the $1g_{9/2}$ level to no more than two nucleons above the minimum occupation for all isotopes. Here, we consider an unrestricted $f_{5}pg_{9}$ model space with the JUN45 [46] and jj44b [47] interactions.
The experimental energy spectra are plotted against the theoretical shell model ones for the even isotopes in Fig. 6 (Appendix B.2). The spectra comparison for 73Ge is provided in [39], alongside a brief summary of the two aforementioned shell model interactions, where the theoretical spectra were found to be inconsistent with another, and not particularly consistent with experiment. In the cases of 70,72,74Ge, the low-lying experimental spectra are a little better reproduced by the JUN45 interaction results, which was fitted to the Ge isotopes in consideration (the jj44b interaction on the other hand was not fitted to these). The shell model spectra for 76Ge are practically identical to one another up to $\sim 1.9$ MeV. An agreement between theoretical and experimental levels of $\sim 200$ keV would be considered satisfactory for such shell model calculations using heavier nuclei, however this agreement exists only for some levels of the even isotopes.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided for the even isotopes in Tables 9 and 10. Much like with 40Ar, the magnetic and electric moment information is not available for the ground state $0^{+}$, however the subsequent non-zero states are used here as a test of the overall accuracy of the shell model code. As before, the values for 73Ge are provided in [39]. For the magnetic dipole moments of the even isotopes, the experimental $2_{1}^{+}$ and $4_{1}^{+}$ states are well reproduced by both shell model interactions, whereas the jj44b value of the $2_{2}^{+}$ state is consistently closer to experiment compared to JUN45. The $2_{1}^{+}$ electric quadrupole moment is quite inconsistent with the experimental values for 70,72,74Ge, where only the jj44b 76Ge value lies within the experimental range.
The individual germanium form factors are plotted (for the $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ responses) in Appendix D, and compared against the SHF and charge distribution form factors. The $M$ response is consistent amongst all calculations, with a maximum difference of $\sim 0.2\%$ at the lower end of the COHERENT $q$ region-of-interest (ROI) for the even isotopes, and $\sim 3\%$ at the higher end for 70,72Ge. For 73Ge these take on the values of $\sim 0.3\%$, and $\sim 20\%$ at the higher end for 73,74,76Ge. For the $\Phi^{\prime\prime}$ channel, the maximum difference is observed between the GCN2850 and jj44b shell model calculations, which is $\sim 80\%$ for 70Ge at both ends of the $q$ spectrum, factor of $\sim 2.4$ for 72Ge, factor of $\sim 2.3$ for 73Ge, factor of $\sim 2.2$ for 74Ge, and $\sim 60\%$ for 76Ge. The difference in the SD $\Sigma^{\prime}$ channel for 73Ge is $\sim 30\%$ at both ends of the $q$ ROI.
127I
The iodine calculations were performed in the model space which includes all proton and neutron orbits in the major shell between magic numbers 50 and 82, with single particle levels $1g_{7/2}$, $2d_{5/2}$, $2d_{3/2}$, $3s_{1/2}$ and $1h_{11/2}$. The work of [35, 38] utilises a nuclear shell model interaction developed by Baldridge and Dalton [48], which we refer to as “B&D” [39]. The valence space restriction employed in [35, 38] involves fixing the occupation number of $1h_{11/2}$ to the minimum allowed nucleon number.
Here, we consider the SN100PN [49] and GCN5082 [45] interactions, and perform two calculations for each. Our shell model calculations are performed in the same model space as the B&D calculations of [35, 38], but with different restrictions. The model space restrictions employed in our work are the same as those in [39], with the four different calculations given the names “SN100PN” and “GCN5082” (which share the same truncation), as well as “Truncated SN100PN” (with a more truncated scheme), and “GCN5082 Expanded” (with an alternative truncation to “GCN5082”). The theoretical energy spectra are plotted against the experimental ones in [39], in addition to the theoretical and experimental electric quadrupole and magnetic dipole moments, with the proton and neutron effective charges and $g$ factors also provided in [39].
The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and charge distribution form factors in Appendix D. The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 30\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 55\%$ and $\sim 50\%$, respectively. For $\Sigma^{\prime}$, the factor differences are $\sim 44$ and $\sim 81$, due to the large difference between the Truncated SN100PN calculation and the remaining calculations. This is consistent with the integrated form factor (IFF) results in [39], which showed a more pronounced difference between the SD Truncated SN100PN IFF results and other calculations. This indicates that a more severe valence space truncation for a shell model interaction has a larger effect compared to a second more comparable truncation.
133Cs
Here, we consider the same valence space and shell model interactions (SN100PN and GCN5082) as 127I above. We have performed four calculations for 133Cs, two of which employ the same truncations as the “SN100PN” and “GCN5082” calculations in Section IV (and are given the same names here). We consider two additional calculations, one each for the SN100PN and GCN5082 interactions, where the neutron $2d_{5/2}$ level now has a minimum of 4 neutrons (as opposed to previously being completely filled), with the remainder of the proton and neutrons levels retaining the previous truncations. This is to explore the effect of a less truncated valence space on the final results. These two additional calculations are termed “SN100PN Expanded” and “GCN5082 Expanded”.
The experimental energy spectra are plotted against the theoretical values for all the above calculations in Fig. 7. Only the SN100PN result reproduces the correct ground state spin and parity of $7/2^{+}$, however all four calculations reproduce the correct spins and parities of the three lowest-lying states within $200$ keV of experiment. For both shell model interactions, using a less truncated valence space causes the spacing of the levels to widen compared to the original calculations, generally pushing the spectrum further up. This is likely due to the opening up of the $2d_{5/2}$ level in these calculations.
The experimental and theoretical magnetic dipole and electric quadrupole moments and transitions are provided in Tables 11 and 12. The proton and neutron effective charges and $g$ factors employed here are the same as those used in [39] for 127I. The experimental magnetic dipole moments are much better reproduced by the shell model calculations compared to the electric quadrupole ones. The experimental electric quadrupole transitions are only reproduced within the uncertainty margins by the GCN5082 calculation, specifically for transitions from the $5/2_{1}^{+}$ and $3/2_{1}^{+}$ states to the ground state. On the other hand, the theoretical magnetic dipole transition values are not consistent with experiment. However, the important test of the shell model wavefunction is the magnetic dipole moment, which is sensitive to the wavefunction of the single state, and to the single nucleon structure of the state, rather than to emerging collective properties, which affect the E2 transition rates and the quadrupole moments. The magnetic dipole transitions are suppressed among the dominant low-excitation configurations, as is evident from their small values, and so become very sensitive to small component in the wavefunction that are not of interest here. Despite the lack of agreement for some observables, the present shell model calculations can be trusted to include the important single-particle configurations in the ground-state wavefunction.
The $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ form factors for the above calculations are compared against the SHF and KN ones in Appendix D. The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 20\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 6\%$ and $\sim 7\%$, respectively. For $\Sigma^{\prime}$, these differences are $\sim 13\%$ and $\sim 18\%$.
@@
Table 4: Our CEvNS cross section predictions in units of 10-40cm2
@@
V Experimental observables
To date, CEvNS has been observed by the COHERENT collaboration using the stopped pion source of neutrinos from the SNS, with both cesium-iodide [50] and argon [2] detector targets. Presently COHERENT have also deployed a germanium detector and measurements are ongoing. There are also experiments aiming to measure CEvNS from reactor neutrinos, including CONUS [51], RICOCHET [52] and NUCLEUS [53]. However, the momentum transfer from reactor neutrinos is very small and thus the form factor has little effect on the cross section. We therefore choose to focus our present analysis on the cesium-iodide, argon and germanium targets of COHERENT and leave reactor based experiments to a future work.
V.1 Flux averaged cross sections
First we calculate the flux averaged cross sections, by performing a weighted average of the cross sections over the three incoming neutrino flavor spectra and the isotopes in the target:
@@ V.1 Flux averaged cross sections
Here the neutrino spectra, $\frac{d\phi}{dE_{\nu}}$, are normalized to unity. For this calculation we use the neutrino spectra provided in [54], which were derived from detailed GEANT4 simulations of the proton beam colliding in the mercury spallation target. The uncertainties in the cross sections are determined from two sources: the form factors and the nucleon charges. We neglect uncertainties in the modelling of the neutrino spectra, which are unavailable, but note that errors in the spectra should affect all of the compared cross sections approximately the same. The form factor uncertainties are found by computing the flux averaged cross section with our two extremal form factors for each model (i.e. the upper and lower limits of the bands shown in Fig. 2). The central value and uncertainty is then taken as ($\sigma_{\rm{max}}+\sigma_{\rm{min}}$)/2 and ($\sigma_{\rm{max}}-\sigma_{\rm{min}}$)/2, respectively. The uncertainty in the nucleon charges is taken from [23]. The total uncertainty is determined by summing these two sources in quadrature. The KN and SHF models are calculated using Eq. (II), while the shell model result uses the full cross section in Eq. (3). The results of this calculation are shown in Fig. 1 and provided numerically in table 4.
Our argon results are all in very close agreement with each other, owing to the fact that argon is an even-even nucleus which doesn’t receive large spin-orbit, and no spin-dependent, contributions. This result is slightly larger than the COHERENT prediction, though with the large uncertainties, is still in reasonable agreement. Our detailed nuclear results do shrink the uncertainty band, though it will be some time before the statistical errors in the experimental results make this difference relevant.
Our cesium-iodide results are in good agreement with the COHERENT prediction. Here we do see a small difference between our shell model results and the SHF calculations. While both cesium and iodine are odd nuclei, the axial term is a positive contribution and so the resulting difference stems from the spin-orbit term.
Our germanium results show a similar trend to the cesium-iodide results, with the shell model results being slightly smaller than the SHF results, which are both with the prior uncertainty derived from charge distributions. Presently there are no COHERENT results to compare with, though they are expected in the near future.
@@
We note that the assumed cross section model will affect the experimentally determined flux averaged cross section through the calculation of the experiment’s sensitivity. Thus as measurements become better, one must be cautious to not interpret deviations of experimental results from predictions which used a different cross section model. For a direct comparison with the observable experimental results we also compute the predicted number of events for COHERENT’s detectors.
V.2 Experimental event counts

### v2

@@ III.1 Evaluating the nuclear responses
where $J_{i}$ is the nuclear ground state angular momentum, and $X=\{M,\,\,\Phi^{\prime\prime}\}$ and $\Sigma^{\prime}$ are nuclear operators whose full form is given in Appendix C.
The nature of these operators can be understood from their long-wavelength ($q\rightarrow 0$) limit (provided in Table 3). The $M$ operator is spin-independent (SI) and adds coherently from all nucleons, giving rise to the standard coherent enhancement of CEvNS. The $\Phi^{\prime\prime}$ term in the $q\rightarrow 0$ limit is proportional to a spin-orbit operator, and is sensitive to nucleon occupations of the spin-orbit partner levels. This partially coherent operator is highest when one spin-orbit partner is filled and the other is empty. The $\Sigma^{\prime}$ operator is spin-dependent (SD), picking out the transverse projection of the spin current. All nuclei have non-zero responses to the $M$ and $\Phi^{\prime\prime}$ operators, while only odd-nuclei with a $J_{i}\neq 0$ ground state have a non-zero $\Sigma^{\prime}$ response.
@@
Using the Wigner-Eckhart theorem the nuclear matrix elements in Eq. (III.1) can be reduced in angular momentum and isospin, giving:
@@
where $\tilde{a}_{|\beta|}=(-1)^{j_{\beta}-m_{j_{\beta}}+1/2-m_{t_{\beta}}}\ a_{|\beta|;-m_{j_{\beta}},-m_{t_{\beta}}}$ and $\otimes$ denotes a tensor product. This basis gives response functions of the form $e^{-2y}p(y)$, where $p(y)$ is a polynomial of the dimensionless parameter $y=(qb/2)^{2}$ and $b={1}/{\sqrt{m_{N}\omega}}\approx\sqrt{41.467/(45A^{-1/3}-25A^{-2/3})}$ fm is the harmonic oscillator length parameter with $\omega$ as the oscillator frequency.
@@
Figure 2: Spin-dependent form factors relevant to CEvNS for the three $J_{i}\neq 0$ targets we consider. The curves denote the different shell model interactions as described in Appendix B.
III.2 Form factors from shell model calculations
Large-scale nuclear shell model calculations were performed using NuShellX [40], which can produce level energies, wave functions and spectroscopic overlaps for a wide range of nuclei. The program’s inputs include the relevant valence space for the nucleus of interest – which constitutes a set of nucleon orbits/levels above a completely filled nuclear core – as well as its corresponding list of two-body interactions. Each set of two-body interactions, also referred to as the ‘interaction’, is fitted to selected experimental nuclear data. The contemporary approach is to derive the interactions from the nucleon-nucleon interaction using effective field theory. The interactions are then fine tuned by fitting a limited number of parameters to a body of nuclear data – usually excitation energies in a range of nuclei relevant to the valence space. The aforementioned OBDMEs are extracted for the valence space levels, with the values dependent on the particular shell model interactions used.
In this work we explore a variety of shell model interactions and truncations for our chosen targets: 40Ar, 70,72,73,74,76Ge, 127I, and 133Cs. Details of the shell model calculations, as well as comparisons with experimental nuclear structure data can be found in Ref. [41] and in Appendix B. The resulting OBDMEs are then used to compute the nuclear responses with the Mathematica package described in [42]. The form factors for each of the nuclear responses $M$, $\Phi^{\prime\prime}$ and $\Sigma^{\prime}$ are plotted as functions of $q$ in Appendix D. Analytic expressions of the form factors for 40Ar, 70,72,73,74,76Ge and 127I are provided in [41], while those for 133Cs are provided in Appendix D. For comparison with our SHF calculations, the form factors are used to derive point-proton, point-neutron and weak radii for all isotopes, and are given in Table 7 of Appendix A.
The shell-model derived weak form factors are shown in Fig. 1, compared to the form factors based on the charge distribution and SHF calculations. The uncertainty band denotes the range of values obtained given the different shell model interactions. The shell model results for each isotope show good agreement with each other (low overall uncertainty) and fall within the large uncertainty band of the form factors derived from the charge distribution. They are also generally in very good agreement with the SHF form factors within the experimentally relevant regions of interest (i.e. the momentum transfer, $q$, which corresponds to observable nuclear recoils in the COHERENT detectors). The axial form factors for each target and interaction are shown in Fig. 2. The germanium and cesium axial form factors exhibit good agreement across the different interactions, while iodine has more disparate results. However, inspecting Eq. (18) and Figs. 1 and 2, we see that the axial term is a factor of (at least) $10^{3}$ smaller than the spin independent term, and is thus a subdominant contribution. This leads us to conclude that the axial form factor uncertainty will be a negligible contribution to the total uncertainty.
@@ III.2 Form factors from shell model calculations
Table 4: Our CEvNS cross section predictions, compared with COHERENT’s predictions and measurements (in units of 10-40cm2)
@@
IV Experimental observables
@@
IV.1 Flux-averaged cross sections
First we calculate the flux-averaged cross sections, by performing a weighted average of the cross sections over the three incoming neutrino flavor spectra and the isotopes in the target:
[…]
@@ IV.1 Flux-averaged cross sections
Here the neutrino spectra, $\frac{d\phi}{dE_{\nu}}$, are normalized to unity. For this calculation we use the neutrino spectra provided in reference [47], which were derived from detailed GEANT4 simulations of the proton beam colliding in the mercury spallation target. The KN and SHF models are calculated using Eq. (II), while the shell model result uses the full cross section in Eq. (18).
The uncertainties in our cross section calculations are determined from two sources: the form factors and the nucleon charges. We neglect uncertainties in the modeling of the neutrino spectra, which are unavailable, but note that errors in the spectra should affect all of the compared cross sections approximately the same. The form factor uncertainties are found by computing the flux averaged cross section with our two extremal form factors for each model (i.e. the upper and lower limits of the bands shown in Fig. 1). The central value and uncertainty are then taken as ($\sigma_{\rm{max}}+\sigma_{\rm{min}}$)/2 and ($\sigma_{\rm{max}}-\sigma_{\rm{min}}$)/2, respectively. The uncertainty in the nuclear weak charge is propagated from the values for the nucleon charges given in Eq. (5) (taken from [25]). The total uncertainty is determined by summing the two sources in quadrature. We find that for our shell model and SHF calculations, the form factor uncertainties are sub-dominant to the nuclear charge uncertainties, with the latter contributing $\sim 0.3\%$.
Figure 3 shows the main result of this work: the flux-averaged cross sections for argon, germanium, and cesium-iodide, calculated using our three approaches. These results are compared with COHERENT’s theoretical prediction (in black) and experimental result (black dashed, with gray area showing the uncertainty band), from [43, 2]. In the case of germanium, our prediction can be used to compare with future COHERENT results, when they become available. The results are also provided numerically in table 4.
The argon cross sections are all in very close agreement with each other, and are slightly larger than the COHERENT prediction, but given the uncertainties, is still in reasonable agreement. Our SHF and shell model results have a much smaller uncertainty band than the COHERENT prediction and their concordance suggests the result is also accurate. The cesium-iodide cross sections are also in good agreement with the COHERENT prediction, falling within the uncertainty band, but again, ours have much smaller uncertainties.
For germanium and cesium-iodide the shell model cross section is smaller than the SHF cross section, with both lying within the uncertainty band of the charge-derived result. There are several potential sources for this difference. Firstly, cesium-133, iodine-127, and germanium-73 are odd nuclei and thus the cross section receives a contribution from the axial current. However, in the previous section we found that the axial nuclear form factor, $F_{A}$, is small. Secondly, our shell model result includes the nuclear spin-orbit response, $\mathcal{F}^{\Phi^{\prime\prime}}$, however this enters the cross section $\propto q^{2}$, and thus, is also small.
To investigate the size of the contributions from the additional nuclear responses, we computed the cross sections using only the shell model derived $F_{M}$ nuclear form factors. For argon, cesium-iodide and germanium, we found that including the axial and spin-orbit terms reduces the flux-averaged cross section by $0.05\%$, $0.05\%$ and $0.08\%$, respectively, well below our margin of error. This shows that the additional nuclear responses have very little effect on the cross section and can thus be neglected at the $<0.1\%$ level of precision.
We therefore conclude that the difference between the SHF and shell model cross sections is due to the $F_{M}$ response. This can be seen visually in Fig. 1, where the argon form factor is the same, while the germanium and iodine (but not cesium) shell model form factors decrease faster than the SHF ones. Note, while Fig. 1 only displays the result for germanium-73, the results are very similar for the other isotopes.
@@
We note that one must be cautious when interpreting deviations of experimental results from predictions. The experimentally determined flux-averaged cross section is typically calculated assuming a theoretical form of the cross section (e.g. through the calculation of an average efficiency factor). Where the energy dependence of the assumed cross section differs from the one used to produce a prediction, a fair comparison cannot be made. While this effect is likely small, for a direct comparison with the experimental results, we now compute the predicted number of events for the given experimental runs.
IV.2 Experimental event counts